OpenStudy (anonymous):

I have just been told that (sin2x)^2=sin^2 2x HOW????

OpenStudy (turingtest):

it's just a difference in notation\[\sin^2x=(\sin x)^2\]

OpenStudy (turingtest):

they mean the same thing; two ways to write it

OpenStudy (turingtest):

note: this is \(not\) the same as\[\sin(x^2)\]

OpenStudy (turingtest):

i.e.\[(\sin x)^2=\sin^2x\neq\sin(x^2)\]

OpenStudy (anonymous):

that kind of goes against what I have learnt about exponents

OpenStudy (anonymous):

I would of said (sin x)^2=sin^2 x^2

OpenStudy (4n1m0s1ty):

but sin() is just a function. its not another variable you can put the exponent on using that rule.

OpenStudy (turingtest):

exactly^ if sin(x) were a variable (sin) times another variable (x) you could distribute but in this case it is a function, and so cannot be simplified that way

OpenStudy (anonymous):

what's really gonna cook you noodle is this notation... \(\sin^{-1}(x^2)\ne [sin(x^2)]^{-1} \)

OpenStudy (4n1m0s1ty):

thats what arcsin() is for